OpenStudy (pricejoin):

Problem Page Solve for w . =+4w3w56 Simplify your answer as much as possible.

KjSaif (kjsaif):

i can't read the question right.....

KjSaif (kjsaif):

w . =+4w3w56 what is this?

KjSaif (kjsaif):

could you send it again :)

OpenStudy (brooke..help00):

w=3wx4wx56? w.=3w4w56? w+3w+4w=56? I don't know the correct equation or I could help.

KjSaif (kjsaif):

yea............

OpenStudy (brooke..help00):

The x's resemble multiplication , sorry for the confusion..

OpenStudy (brooke..help00):

Ok, it says to solve for w. (Which I am guessing the w . is the end of the sentence. If that is correct it looks like the equation would be 3w(Times)4w (times 56)

OpenStudy (brooke..help00):

\[3w \times 4w \times 56\]

OpenStudy (brooke..help00):

\[12w \times 56\]

OpenStudy (brooke..help00):

\[w = 4.6 or 4.5?\]

OpenStudy (mathmale):

Kindly repost your question. Your "+4w3w56" is too ambiguous to make sense. Do you mean "exponentiation" when you type w3? If so, that should be w^3, not w3. Use parentheses, if necessary, to indicate order of operations intended.

OpenStudy (dallascowboys88):

You Have To Use The Method PEMDAS.

OpenStudy (dallascowboys88):

Parenthesis,Equalaties,Multiply,Divide,Add,Subtract.In That Order.

OpenStudy (mathmale):

If there's exponentiation here, then yes, you have to do the exponentiation first, according to "PEMDAS." If it's all multiplication, then the order in which you do the mult. doesn't matter. Again, kindly clarify this problem to remove the ambiguities.

OpenStudy (brooke..help00):

And exponents not equalities

OpenStudy (mathmale):

Please repost your question. Indicate any exponentiation with " ^ ".

OpenStudy (brooke..help00):

@pricejoin We all would love to help you! We jut need you to repost your question so we can understand it :)